what is the probability of picking two cards of the same suit in a deck of 52?
Answer by stanbon(75887)

what is the probability of picking two cards of the same suit in a deck of 52?
---------------
Pick a suit: 4 ways
Pick two cards from that suit: 13C2 = 13*12/1*2 = 13*6 = 78
---------------------
Unrestricted number of pairs of cards: 52C2 = 52*51/1*2 = 26*51 = 1326
----------------------
P(pick 2 cards of the same suit) = [4*78]/1326 = 0.23529...
================
Cheers,
Stan H.